The plot shows the position versus time of a mass of 1.2 kg on a spring. What is the spring constant?

dynamically generated plot

Mass on Spring

Can you show work please?
Answer #1

from the graph

time period of Oscillations, T = 20/2

= 10 s

so, angular frequency, w = 2*pi/T

= 2*pi/10

= 0.6283 rad/s

given m = 1.2 kg
let k is the spring constant

we know, w = sqrt(k/m)

w^2 = k/m

==> k = m*w^2

= 1.2*0.6283^2

= 0.474 N/m <<<<<<<------------------Answer